# Single-site dissipation stabilizes a superconducting nonequilibrium steady state in a strongly correlated lattice

### v1

@@ V.2 Disorder that suppresses or destroys the $\eta$-pair ODLRO plateau
Figure 11: Pair-breaking single-particle loss with rate $\kappa$ (normalized as $\kappa/\gamma$) strongly suppresses the $\eta$-paired NESS. (a) $|\Phi_{m}|_{\mathrm{ss}}$ and (b) the long-distance plateau $|C_{m}(r\!\to\!N/2)|_{\mathrm{ss}}$ versus $\kappa/\gamma$. Dashed lines indicate the projected benchmarks $1/2$ and $1/4$ at $\kappa=0$. Parameters: 1D chain with PBC, $t=\gamma$, $U=6/\gamma$, $\theta=\pi/2$, $N=5$, and dissipation applied on site 22. Turning on $\kappa$ induces a leak-and-scramble mechanism: $|C_{m}(r\!\to\!N/2)|_{\mathrm{ss}}$ collapses rapidly toward zero, while $|\Phi_{m}|_{\mathrm{ss}}$ is reduced more gradually due to short-time local repumping but remains bounded by the continuous pair breaking.
@@
Finally, we consider additional uncontrolled dissipators that are not aligned with the $\eta$-pumping protocol, the most important example being single-particle loss,
$L_{i\sigma}^{\mathrm{loss}}$ $=$ $\sqrt{\kappa}\,c_{i\sigma}\qquad$ (54)
$(\text{and similarly dephasing }L_{i\sigma}^{\mathrm{dep}}$ $=$ $\sqrt{\kappa_{\phi}}\,n_{i\sigma}).$ (55)
@@
Such channels are qualitatively different from the disorder terms discussed above: they act as genuine pair-breaking processes. In the local Fock basis, $c_{i\sigma}$ converts holons and doublons into singly occupied states,
@@
$c_{i\sigma}:\ |0\rangle_{i}\mapsto 0,\quad|\uparrow\downarrow\rangle_{i}\mapsto|\bar{\sigma}\rangle_{i},\qquad c_{i\sigma}^{\dagger}:\ |0\rangle_{i}\mapsto|\sigma\rangle_{i},$ (56)
@@
thereby opening real dissipative leakage channels from the holon–doublon doublet $\mathrm{span}\{|0\rangle_{i},|\uparrow\downarrow\rangle_{i}\}$ into the singlon sector $\{|\uparrow\rangle_{i},|\downarrow\rangle_{i}\}$. In the $\eta$ language this is precisely the destruction of the locally pumped coherence: the designed jump $L_{j}\propto\eta_{j}^{-}e^{i\frac{\pi}{2}\eta_{j}^{x}}$ can only rephase and repump within the $\eta_{i}=1/2$ doublet, whereas single-particle loss continuously removes weight from that sector and scrambles the relative phase between $|0\rangle_{i}$ and $|\uparrow\downarrow\rangle_{i}$.
@@
Numerically, this leak-and-scramble channel is strongly detrimental to $\eta$-pair coherence. As soon as the loss rate $\kappa$ is nonzero, the long-distance correlator $|C_{m}(r\!\to\!N/2)|_{\mathrm{ss}}$ is rapidly suppressed toward zero, signaling a prompt destruction of ODLRO; see Fig. 11. The one-point amplitude $|\Phi_{m}|_{\mathrm{ss}}$ is also reduced, typically more gradually, because local rotated pumping can transiently repolarize the $\eta$ pseudospin and sustain short-range coherence, but it cannot compensate an ongoing pair-breaking process at long times. In this sense, uncontrolled loss imposes a stringent upper bound on the steady-state ODLRO plateau and, for fixed pumping strength $\gamma$, sets the achievable quality (and effective lifetime) of the $\eta$-paired nonequilibrium steady state.
In summary, the disorder response supports a symmetry-based robustness principle: perturbations that are diagonal (or effectively longitudinal) in the local HD doublet mainly renormalize relaxation pathways and leave the ODLRO plateau close to the projected value, whereas perturbations that (i) compete directly with the dissipative axis selection (e.g., transverse fields or angle disorder) or (ii) open real leakage/pair-breaking channels (e.g., single-particle loss) suppress and ultimately destroy the nonequilibrium superconducting order.
@@ VI Conclusion and Outlook
We have demonstrated that superconducting-like order can be stabilized as a nonequilibrium attractor in a strongly correlated lattice system. For the particle–hole symmetric Hubbard chain, we introduced a rotated quantum-jump protocol based on a locally rotated $\eta$-pair lowering operator and showed that the Lindblad dynamics converges from the vacuum to a NESS with $\eta$-pair ODLRO.
The mechanism underlying our protocol can be understood through three essential components. First, local dark-state engineering: the rotated quantum jump operator reshapes the local dark-state manifold and imprints a well-defined $\eta$-pseudospin phase, thereby establishing a robust on-site coherence between holon and doublon configurations. Second, selective pumping governed by the representation structure of the dissipator and dynamical separation of timescales: since the local $\eta$ operators act nontrivially only within the holon–doublon doublet while annihilating singlon states, the dissipative dynamics does not directly generate singly occupied configurations. Any singlon admixture originates solely from coherent hopping processes and remains perturbative in the regime where fast dynamical components can be systematically eliminated via a controlled Schur-complement expansion. Third, coherence propagation mediated by coherent hopping together with Liouvillian selection: the imprinted pseudospin phase spreads across the lattice through coherent dynamics, while the projected Liouvillian dynamics selects the $\eta$-paired subspace as an attractive manifold and opens a dissipative gap that governs the relaxation timescale toward the steady state. A notable consequence of this mechanism is a local-to-global preparation principle: dissipative action confined to a single lattice site can induce macroscopic ODLRO throughout the interacting many-body system.
We further established that this $\eta$-paired NESS is robust against broad classes of static disorder, which mainly renormalize transient dynamics and relaxation rates, and we identified specific perturbations that directly compete with dissipative axis selection (or introduce pair breaking) and thus suppress long-range order. Looking forward, several directions are immediately suggested. First, it would be valuable to develop analytic scaling descriptions of the dissipative gap and relaxation with system size, driving geometry, and disorder strength. Second, the mechanism is not tied to one dimension: since $\eta$-SU(2) structure and the staggered $\eta$-pair generators are naturally compatible with any bipartite lattice, the protocol should extend directly to arbitrary bipartite geometries (and, in particular, to higher-dimensional lattices) with appropriate choices of driven sites. Third, spatially patterned or time-dependent rotations/jumps offer a natural route to imprint phase textures, domain structures, and supercurrents in a controlled manner. More broadly, our results illustrate a local-to-global preparation principle: minimal, local dissipation engineering can autonomously stabilize superconducting order as an attractor of strongly correlated open-system dynamics.

### v2

@@ V.2 Disorder that suppresses or destroys the $\eta$-pair ODLRO plateau
Figure 11: Pair-breaking single-particle loss with rate $\kappa$ (normalized as $\kappa/\gamma$) strongly suppresses the $\eta$-paired NESS. (a) $|\Phi_{m}|_{\mathrm{ss}}$ and (b) the long-distance plateau $|C_{m}(r\!\to\!N/2)|_{\mathrm{ss}}$ versus $\kappa/\gamma$. Dashed lines indicate the projected benchmarks $1/2$ and $1/4$ at $\kappa=0$. Parameters: 1D chain with PBC, $t=\gamma$, $U=6\gamma$, $\theta=\pi/2$, $N=5$, and dissipation applied on site 22. Turning on $\kappa$ induces a leak-and-scramble mechanism: $|C_{m}(r\!\to\!N/2)|_{\mathrm{ss}}$ collapses rapidly toward zero, while $|\Phi_{m}|_{\mathrm{ss}}$ is reduced more gradually due to short-time local repumping but remains bounded by the continuous pair breaking.
@@
Finally, we consider additional uncontrolled dissipators that are not aligned with the $\eta$-pumping protocol, the most prominent example being single-particle loss,
$L_{i\sigma}^{\mathrm{loss}}$ $=$ $\sqrt{\kappa}\,c_{i\sigma},$ (53)
$L_{i\sigma}^{\mathrm{dep}}$ $=$ $\sqrt{\kappa_{\phi}}\,n_{i\sigma},$ (54)
@@
where the latter represents local dephasing. These channels are qualitatively distinct from the disorder terms previously discussed because they act as genuine pair-breaking processes. In the local Fock basis, $c_{i\sigma}$ converts holons and doublons into singly occupied states,
@@
$c_{i\sigma}:|0\rangle_{i}\mapsto 0,\quad|\uparrow\downarrow\rangle_{i}\mapsto|\bar{\sigma}\rangle_{i},\qquad c_{i\sigma}^{\dagger}:|0\rangle_{i}\mapsto|\sigma\rangle_{i}.$ (55)
@@
Such processes open real dissipative leakage channels from the holon–doublon doublet into the singlon sector. In the $\eta$-pseudospin language, this corresponds to the irreversible destruction of local coherence. While the designed jump $L_{j}$ can rephase and repump within the $\eta_{i}=1/2$ doublet, single-particle loss continuously depletes the weight of that sector and scrambles the relative phase between $|0\rangle_{i}$ and $|\uparrow\downarrow\rangle_{i}$.
@@
Numerically, this leak-and-scramble mechanism is highly detrimental to $\eta$-pair coherence. As soon as the loss rate $\kappa$ becomes nonzero, the long-distance correlator $|C_{m}(r\to N/2)|_{\mathrm{ss}}$ is rapidly suppressed, signaling the prompt destruction of ODLRO as shown in Fig. 11. The one-point amplitude $|\Phi_{m}|_{\mathrm{ss}}$ is also reduced, albeit typically more gradually, as the local rotated pumping can transiently repolarize the $\eta$-pseudospin and maintain short-range coherence. However, the pumping cannot compensate for an ongoing pair-breaking process at long times. Consequently, uncontrolled loss imposes a stringent upper bound on the steady-state ODLRO plateau and, for a fixed pumping strength $\gamma$, determines the achievable quality and effective lifetime of the $\eta$-paired NESS.
In summary, the response to disorder can be organized according to a symmetry-based robustness principle characterized by two complementary failure modes. Perturbations that act diagonally, or effectively longitudinally, within the local holon–doublon doublet typically do not compete with the dissipative axis selection. As long as off-manifold excursions remain perturbative, these terms primarily renormalize relaxation pathways while leaving the ODLRO plateau near its ideal projected value. This category includes bond inhomogeneity of bipartite AABB hopping, which fails to generate coherent dynamics within the $\eta$-multiplet at leading order. By contrast, the nonequilibrium superconducting order is suppressed and eventually destroyed when perturbations either tilt the local dissipative axis within the reduced manifold, such as in the case of transverse fields and angle disorder, or cause the off-manifold resolvent to become large. The latter case opens real leakage and pair-breaking channels, as observed with strong potential disorder and single-particle loss.
@@ VI Conclusion and Outlook
In this work, we have demonstrated that superconducting-like order can be stabilized as a nonequilibrium attractor in a strongly correlated lattice system. For the particle–hole symmetric Hubbard chain, we introduced a rotated quantum-jump protocol based on a locally rotated $\eta$-pair lowering operator. We showed that the resulting Lindblad dynamics pumps the system from the vacuum into a nonequilibrium steady state (NESS) characterized by $\eta$-pair off-diagonal long-range order (ODLRO).
The mechanism can be summarized as follows. The rotated jump reshapes the local dark-state structure and pins a definite $\eta$-pseudospin phase, thereby selecting coherent holon–doublon superpositions as local building blocks. In the strong-coupling regime, singlon configurations are not populated by the dissipator to leading order and enter only through hopping-induced virtual excursions, which remain perturbative and can be eliminated using a controlled Schur-complement expansion. Coherent hopping then propagates the imprinted pseudospin phase across the lattice, while the projected Liouvillian dynamics selects the $\eta$-paired manifold as an attractive sector and opens a dissipative gap that sets the relaxation timescale. A notable consequence is a local-to-global preparation principle: dissipation applied on a single site can induce macroscopic ODLRO throughout the interacting system.
We further found that the $\eta$-paired NESS is robust against broad classes of static disorder, which mainly renormalize transients and relaxation rates, and we identified perturbations that suppress long-range order by competing with dissipative axis selection or by opening genuine leakage and pair-breaking channels.
Looking forward, it will be valuable to develop analytic scaling theories for the dissipative gap and relaxation as functions of system size, driving geometry, and disorder strength. It is also natural to explore extensions to higher-dimensional bipartite lattices and to investigate spatially patterned or time-dependent rotations and jumps as tools to imprint phase textures, domains, and currents in a controlled and autonomous manner. More broadly, our results highlight minimal local dissipation as a practical route to stabilizing correlated order in driven open systems.